# The Cosmic Evolution and Spatial Distribution of Multiphase Gas associated with QSOs

### v1

@@ Abstract
We investigate the multi-phase gas surrounding QSOs traced by 33 absorption lines (e.g., Ly, C iv, Fe ii, Mg ii, etc.) in the stacked spectra of background sources, using the early data release from the Dark Energy Spectroscopic Instrument. Our analysis reveals that the equivalent width () of metal absorption lines decreases with increasing redshift, following an overall trend described by . This indicates cosmic enrichment of the gas associated with QSOs, with a more pronounced enrichment observed at larger scales (a few Mpc), reflecting a rapid metal enrichment of cosmic web. Additionally, the of these absorption lines decreases with distance () from QSOs, which can be effectively characterized by a two-halo model. Compared to the projected two point correlation function of galaxies at similar redshifts, low-ionization ions exhibit similar clustering scales, while high-ionization ions show a significantly more extended spatial distribution. We also find that increases towards lower redshifts and larger scales, which can be attributed to evolving star formation histories and/or changes in initial mass function for galaxies. By leveraging multiple absorption tracers, we conduct the first comprehensive investigation of diffuse, multiphase gas from the circumgalactic medium to cosmological scales, offering new insights into baryon cycles and the transport of metals throughout cosmic time.
@@ 1 Introduction
Metal-enriched absorption is widely believed to result from large-scale galactic outflows driven by stellar winds, supernovae (SNe), active galactic nuclei (AGN), and the accretion of gas tidally stripped during mergers from star-forming galactic discs (e.g., Oppenheimer & Davé, 2006; Dalla Vecchia & Schaye, 2008; Suresh et al., 2015). Compared to massive galaxies, dwarf galaxies offer more opportunities for metals to be transferred to the CGM and even the Intergalactic Medium (IGM) through these outflows (Muratov et al., 2017; Ayromlou et al., 2023). Observational evidence for such outflows is well-established through the detection of blueshifted Mg ii and Fe ii absorption lines observed down-the-barrel in galactic spectra, which trace kinematically disturbed gas moving at velocities of 100 to 1000 km/s (e.g. Weiner et al., 2009; Erb et al., 2012; Kornei et al., 2012; Martin et al., 2012; Rubin et al., 2014). These feedback mechanisms facilitate the bulk transport of metals from disk to halo scales and establish a dynamic cycling process, redistributing enriched material throughout the CGM and the cosmic web (Bordoloi et al., 2011; Prochaska et al., 2011; Rubin et al., 2014).
Despite the ability of outflow mechanisms to explain the origin of metals in the gas surrounding galaxies (Shin et al., 2021), substantial uncertainties remain regarding the spatial distribution (e.g., Péroux et al., 2020; Weng et al., 2024) and temporal evolution (e.g., Dey et al., 2015; Lan, 2020; Wu et al., 2025) of these metals once expelled from the galactic disk. This process involves complex interactions within the multiphase medium, such as outflow deceleration, gas mixing, reaccretion, and interactions with cosmic filaments (Tumlinson et al., 2017; Faucher-Giguère & Oh, 2023; Saeedzadeh et al., 2023). The retention time, diffusion extent, and chemical state of metals in the gas surrounding galaxies are heavily influenced by the physical conditions of the transport paths they transverse (Maiolino & Mannucci, 2019; Fumagalli, 2024).
The Eris simulation indicates that hot metals () preferentially propagate into the CGM via high-velocity outflows, establishing a radially declining temperature gradient (Shen et al., 2012). Cool photoionized gas () is primarily traced by the C ii and Si ii absorption lines, which exhibit a sharply decreasing covering factor beyond . In contrast, high-ionization species such as O vi (e.g., Sanchez et al., 2019; Li et al., 2021) and C iv maintain a substantial covering factor beyond , with O vi-enriched halos extending to (Shen et al., 2013). These results align with observations of star-forming galaxies at low redshift by Tumlinson et al. (2011) and Prochaska et al. (2011).
@@ 2.1 DESI EDR
The DESI (DESI collaboration et al., 2016a, b) is the largest multi-object spectrograph ever constructed, designed to conduct a large-scale spectroscopic survey to probe the nature of dark energy (DESI collaboration et al., 2022). Installed on the 4-meter Mayall Telescope at Kitt Peak National Observatory, DESI enables a survey of unprecedented scale, depth, and efficiency (Levi et al., 2019), mapping tens of millions of galaxies and quasars to study cosmic expansion.
@@
Galaxies in DESI EDR can be categorized into three classes (Myers et al., 2023): Emission Line Galaxies (Raichoor et al., 2020, 2023), Luminous Red Galaxies (Zhou et al., 2020, 2023), and Bright Galaxy Survey (BGS) (Ruiz-Macias et al., 2020; Hahn et al., 2023). We utilize the SPECTYPE from the zcatalog222https://data.desi.lbl.gov/public/edr/spectro/redux/fuji/zcatalog/zall-tilecumulative-fuji.fits of the DESI project to distinguish between QSOs (Yèche et al., 2020; Chaussidon et al., 2023) and galaxies. The redshift estimation algorithm for galaxies in this catalog is known as redrock333https://github.com/desihub/redrock. The stellar masses of galaxies are obtained from the value-added catalog 444https://data.desi.lbl.gov/public/edr/vac/edr/stellar-mass-emline/v1.0/edr_galaxy_stellarmass_lineinfo_v1.0.fits (Zou et al., 2019), which estimates these masses using the CIGALE SED fitting code (Boquien et al., 2019).
@@ 2.2 Spectral stacking method
The median-filtered spectra were then co-added onto a wavelength grid with an interval of 0.2 Å, where the median value was adopted as the flux of the stacked spectrum. We employed the bootstrap resampling method to estimate the uncertainty of the normalized flux at each wavelength pixel, using 100 realizations. Finally, we performed Gaussian fitting of the absorption features using the emcee (Foreman-Mackey et al., 2013) and integrated the flux within the absorption troughs to derive the equivalent width. This stacking approach is widely adopted as an unbiased method due to its flexibility and reliability (e.g., Bordoloi et al., 2011; Zhu & Ménard, 2013b, a; Zhu et al., 2014; Bordoloi et al., 2014).
We present a stacked spectrum with a foreground QSO (, ) in Figure 1. The spectrum is divided into two segments (top and bottom panels) within the range of 1100 2900 Å and an appropriate y-axis range is selected to simultaneously display absorption lines with varying intensities. As shown, after stacking hundreds of thousands of spectra, the fluctuations and errors in the stacked spectrum are minimized, enabling the detection of very faint absorption troughs. The complete spectral data can be found on our website 666https://drive.google.com/drive/folders/1pjHTLGIQzEu6gGoHEx8bTVZqFiB81Uu7?usp=sharing.
@@ 3 RESULTS
We present the (in mÅ) of all detected absorption lines using QSOs as foreground samples in Table LABEL:tab1. The data are organized into four redshift bins: , , , and , along with three impact parameter bins: , , and . The average impact parameters for three bins are 0.8 Mpc, 2.5 Mpc, and 5.1 Mpc, respectively. A blank entry indicates that the absorption line wavelength exceeds the limits of the co-added spectrum, while a dash denotes cases where the is too low to fit the absorption line and measure the . Additionally, all measurements with are marked with an asterisk ( * ) in Table LABEL:tab1. Due to space limitations, errors for are not included here; however, a complete table with errors is available on our website 777https://drive.google.com/drive/folders/1pjHTLGIQzEu6gGoHEx8bTVZqFiB81Uu7?usp=sharing. In this study, we select QSOs as foreground pairs for the main sample, as they provide a broad redshift distribution of foreground objects, facilitating a more comprehensive examination of evolutionary effects. The equivalent width results for ELGs and LRGs are included in Appendix A, and the corresponding data, including the stacked spectra and absorption measurements, can be accessed on the same website. The impact parameter bins, ranging from 0.02 to 6.5 Mpc, encompass the CGM and the gas that connects to the large-scale structure of the universe, making them suitable for conducting a thorough survey of the gas associated with QSOs.
@@ 3.1 Evolution of with redshift and enrichment rates
We plot the data with a from Table LABEL:tab1 in Figure 2. This criterion is consistently applied to all figures within this paper. To ensure the robustness of the plotted points, we require that each absorption line has at least two equivalent width measurements within four redshift bins. The upper panels display the equivalent widths () using different marker shapes to represent each element, with colors indicating their respective rest wavelengths. This marking scheme is consistent across all figures in this work, except for Figure 6, which focuses on the ratio of two lines. The lower panels of Figure 2 present the normalized equivalent widths () as a function of .
A consistent negative correlation is observed across all ions, where systematically decreases as redshift increases. The lower panels illustrate this trend more distinctly. The normalization of accounts for differences in the initial values, facilitating clearer comparisons of trends across various radial bins and redshifts. The observed decrease in with redshift is evident in all radial bins with slopes of (0.8 Mpc), (2.5 Mpc), and (5.1 Mpc). Notably, the dependence on redshift is stronger at an impact parameter of 5.1 Mpc, () compared to smaller scales.
Galactic feedback is crucial in shaping the multiphase structure of the CGM at distances less than , whereas regions beyond this are more influenced by cosmological effects, such as non-spherical accretion of intergalactic gas (Fielding et al., 2020; Li et al., 2021). This trend suggests that over cosmic time, metals produced and released by both star formation processes and Active Galactic Nuclei (AGN) feedback contribute to the overall enrichment of metals in the universe (e.g., Moll et al., 2007; Germain et al., 2009; Huang et al., 2023; Villar Martín et al., 2024). Initially, this enrichment occurs in the gas closest to QSOs (or galaxies), and as the inner regions approach a state of near-complete metal enrichment, these metals gradually propagate outward into more distant regions. Consequently, in the later universe, we are more likely to observe enrichment processes taking place in the outer regions. This dynamic also accounts for the faster enrichment at larger , as there is a delay in the propagation of metal enrichment from small to large scales associated with galaxies.
In Figure 3, we present the fitted slopes for each absorption line in the vs. plot (as shown in Figure 2). The evolution of different elements and ions shows significant variability. Notably, ions with rest wavelengths shorter than 1800 Å exhibit a stronger dependence on redshift, which may primarily result from differing ionization states of these absorption lines rather than purely cosmic evolution of metals. Conversely, among all the absorption lines, Mgii shows the weakest evolution, indicating little to no change over time, which aligns with trends noted by Lan (2020), Anand et al. (2021) and Chen et al. (2025). The overall structure of Mg ii-bearing halos was established early in the galaxy assembly process and have evolved relatively little since then (Matejek & Simcoe, 2012).
@@
We also perform the linear fittings to the profiles of these absorption lines as shown in the right panel of Figure 4. The results indicates that the absorption intensity decreases with , which aligns with previous findings (e.g., Zhu & Ménard, 2013a; Lan & Mo, 2018; Méndez-Hernández et al., 2022; Wu et al., 2025). Mg ii absorption here is slightly lower than Zhu et al. (2014) for LRGs at . This discrepancy primarily arises from the higher redshift of the foreground QSOs () in our study. Ly demonstrates a stronger absorption as predicted by Shen et al. (2013) and remains relatively stable on large scales.
The fitted slopes from the linear fittings for each ion are displayed in Figure 5. The observed differences in could be attributed to the varying physical processes that affect the distribution of each ion in the gas surrounding galaxies. For example, high-ionization species such as C iv and Si iv are typically found in hotter, more ionized regions. In contrast, low-ionization species such as Mg ii and Fe ii are generally located in colder, and denser regions. High-ionization species tend to maintain a more extended distribution (with smaller ), as corroborated by both simulations (Shen et al., 2013; Sanchez et al., 2019; Li et al., 2021; Carr et al., 2025) and observational studies (Tumlinson et al., 2011; Prochaska et al., 2011; Méndez-Hernández et al., 2022). Conversely, low-ionization species Mg ii and Fe ii show stronger dependence on impact factor, with values of (Fe ii ), (Fe ii ), (Fe ii ), (Fe ii ), and (Mg ii ). Additionally, in Figure 5, we present the slope of the projected two-point correlation function for galaxies within the 1–5 Mpc range, marked by different colors: Zehavi et al. (2004) for (blue), Nuza et al. (2013) for (green), and Durkalec et al. (2018) for (orange).
@@
The ratio is shown as a function of (left panel) and the impact parameter (right panel) in Figure 6. Different colors correspond to various Fe ii absorption lines indicated in the legend. The results indicate that as the redshift increases, the ratio decreases following a trend proportional to , consistent with Dey et al. (2015). Conversely, as increases, the ratio shows a slight increase. This suggests that Mg ii is likely more concentrated in the inner regions and evolves more slowly with redshift.
Mg, being an -element, is predominantly produced by core-collapse supernovae (CCSN), which are the result of the star explosions. In contrast, iron (Fe) is primarily generated by Type Ia supernovae (SNIa), which originate from the explosion of accreting white dwarfs. The ratio of these elements should reflect the relative frequencies of CCSN and SNIa, as well as the time lag between their productions. In studying the stellar absorption lines, Conroy et al. (2014) found significant enhancements in massive elliptical galaxies formed in the early universe, attributed to the intense bursts of star formation that occurred at high redshift (Li et al., 2025).
@@
In contrast, Fe is predominantly produced by SNIa, where the progenitor stars are white dwarfs in binary systems. These white dwarfs undergo accretion or merger processes that typically take about 1 Gyr (Matteucci & Recchi, 2001; Maoz et al., 2010; Horiuchi & Beacom, 2010). This delay decouples Fe release from the star formation activities associated with dense gas clouds, allowing Fe to be transported to more distant regions following supernova explosion. This may explain the increase with . Additionally, on small scales, Fe may be partially adsorb onto dust particles (Konstantopoulou et al., 2022), reducing its observed abundance in ionized gas. As the impact parameter increases, the dust density decreases, causing Fe to desorb from the dust and re-enter the gas phase, which may contribute to an increase in the [Fe/Mg] ratio.

### v2

@@ Abstract
We investigate the multi-phase gas surrounding QSOs traced by 33 absorption lines (e.g., Ly, C iv, Fe ii, Mg ii, etc.) in the stacked spectra of background sources, using the early data release from the Dark Energy Spectroscopic Instrument. Our analysis reveals that the equivalent width () of metal absorption lines decreases with increasing redshift, following an overall trend described by . Different species that trace multi-phases of QSO-associated gas exhibit distinct evolutionary patterns. Additionally, the of these absorption lines decreases with distance () from QSOs, which can be effectively characterized by a two-halo model. Compared to the projected two point correlation function of galaxies at similar redshifts, low-ionization ions exhibit similar clustering scales, while high-ionization ions show a significantly more extended spatial distribution. We also find that increases towards lower redshifts, which can be attributed to evolving star formation histories and/or changes in initial mass function for galaxies. By leveraging multiple absorption tracers, we conduct the first comprehensive investigation of diffuse, multiphase gas from the circumgalactic medium to cosmological scales, offering new insights into baryon cycles and the transport of metals throughout cosmic time.
@@ 1 Introduction
Metal-enriched absorption is widely believed to result from large-scale galactic outflows driven by stellar winds, supernovae (SNe), active galactic nuclei (AGN), and the accretion of gas tidally stripped during mergers from star-forming galactic discs (e.g., Oppenheimer & Davé, 2006; Dalla Vecchia & Schaye, 2008; Suresh et al., 2015). Compared to massive galaxies, dwarf galaxies offer more opportunities for metals to be transferred to the CGM and even the Intergalactic Medium (IGM) through these outflows (Muratov et al., 2017; Ayromlou et al., 2023). Observational evidence for such outflows is well-established through the detection of blueshifted Mg ii and Fe ii absorption lines observed down-the-barrel in galactic spectra, which trace kinematically disturbed gas moving at velocities of 100 to 1000 km/s (e.g. Weiner et al., 2009; Erb et al., 2012; Kornei et al., 2012; Martin et al., 2012; Rubin et al., 2014). These feedback mechanisms facilitate the bulk transport of metals from disk to halo scales and establish a dynamic cycling process, redistributing enriched material throughout the CGM and the cosmic web (Bordoloi et al., 2011; Prochaska et al., 2011a; Rubin et al., 2014).
Despite the ability of outflow mechanisms to explain the origin of metals in the gas surrounding galaxies (Shin et al., 2021), substantial uncertainties remain regarding the spatial distribution (e.g., Péroux et al., 2020; Weng et al., 2024) and temporal evolution (e.g., Dey et al., 2015; Lan, 2020) of these metals once expelled from the galactic disk. This process involves complex interactions within the multiphase medium, such as outflow deceleration, gas mixing, reaccretion, and interactions with cosmic filaments (Tumlinson et al., 2017; Faucher-Giguère & Oh, 2023; Saeedzadeh et al., 2023). The retention time, diffusion extent, and chemical state of metals in the gas surrounding galaxies are heavily influenced by the physical conditions of the transport paths they transverse (Maiolino & Mannucci, 2019; Fumagalli, 2024).
@@ 2.1 DESI EDR
The DESI (DESI Collaboration et al., 2016a, b) is the largest multi-object spectrograph ever constructed, designed to conduct a large-scale spectroscopic survey to probe the nature of dark energy (DESI Collaboration et al., 2019, 2022). Installed on the 4-meter Mayall Telescope at Kitt Peak National Observatory, DESI enables a survey of unprecedented scale, depth, and efficiency (Levi et al., 2019), mapping tens of millions of galaxies and quasars to study cosmic expansion.
@@
Galaxies in DESI EDR can be categorized into three classes (Myers et al., 2023): Emission Line Galaxies (Raichoor et al., 2020, 2023), Luminous Red Galaxies (Zhou et al., 2020, 2023), and Bright Galaxy Survey (BGS) (Ruiz-Macias et al., 2020; Hahn et al., 2023). We utilize the SPECTYPE from the zcatalog222https://data.desi.lbl.gov/public/edr/spectro/redux/fuji/zcatalog/zall-tilecumulative-fuji.fits of the DESI project to distinguish between QSOs (Yèche et al., 2020; Chaussidon et al., 2023) and galaxies. The redshift estimation algorithm for galaxies in this catalog is known as redrock333https://github.com/desihub/redrock. Here, we apply a selection of ZWARN or to ensure the reliability of the redshift measurements in DESI. The stellar masses of galaxies are obtained from the value-added catalog 444https://data.desi.lbl.gov/public/edr/vac/edr/stellar-mass-emline/v1.0/edr_galaxy_stellarmass_lineinfo_v1.0.fits (Zou et al., 2019), which estimates these masses using the CIGALE SED fitting code (Boquien et al., 2019).
@@ 2.2 Spectral stacking method
The median-filtered spectra were then co-added onto a wavelength grid with an interval of 0.2 Å, where the median value was adopted as the flux of the stacked spectrum. We employed the bootstrap resampling method to estimate the uncertainty of the normalized flux at each wavelength pixel, using 100 realizations. We compared the errors from 100 and 1000 bootstrap realizations and confirmed that the 100 realizations are sufficiently robustness for this work. Finally, we performed Gaussian fitting of the absorption features using the emcee (Foreman-Mackey et al., 2013) and integrated the flux within the absorption troughs to derive the equivalent width. This stacking approach is widely adopted as an unbiased method due to its flexibility and reliability (e.g., Bordoloi et al., 2011; Zhu & Ménard, 2013b, a; Zhu et al., 2014; Bordoloi et al., 2014). Due to the broad wavelength range covered in this work, it is unavoidable that some absorption lines fall within the Ly forest region of the background QSO. The potential impact of this mixed stacking is discussed in detail in Appendix B. As shown, absorption line measurements are nearly unaffected by this effect.
We present a stacked spectrum with a foreground QSO (, ) in Figure 1. The spectrum is divided into two segments (top and bottom panels) within the range of 1100 2900 Å and an appropriate y-axis range is selected to simultaneously display absorption lines with varying intensities. As shown, after stacking hundreds of thousands of spectra, the fluctuations and errors in the stacked spectrum are minimized, enabling the detection of very faint absorption troughs. The complete spectral data are available in our publicly released https://doi.org/10.5281/zenodo.15617674 (catalog dataset) 666https://doi.org/10.5281/zenodo.15617674.
@@ 3 RESULTS
We present the (in mÅ) of all detected absorption lines using QSOs as foreground samples in Table LABEL:tab1. The data are organized into four redshift bins: , , , and , along with three impact parameter bins: , , and . The average impact parameters for three bins are 0.8 Mpc, 2.5 Mpc, and 5.1 Mpc, respectively. A blank entry indicates that the absorption line wavelength exceeds the limits of the co-added spectrum, while a dash denotes cases where the is too low to fit the absorption line and measure the .
Additionally, an asterisk ( * ) in Table LABEL:tab1 indicates either measurements with or absorption features identified as narrow, jitter-induced depressions misinterpreted by Gaussian fitting. Gaussian fitting may occasionally capture narrow spurious absorption features caused by spectral fluctuations (e.g., Fe ii 2249 and Fe ii 2260 in Figure 1). To mitigate this, we impose a criterion that the equivalent width of a line must exceed three times the square root of the summed squared error within a 5 Å window centered on the line. This window is broad enough to suppress spurious features from narrow fluctuations while retaining genuine absorption lines. Nonetheless, final visual inspection remains essential to ensure the reliability of the fits. Measurements of absorption lines that lie too close to the blue end of the stacked spectrum, where spectral coverage becomes insufficient, are also struck through with a horizontal line and excluded from the analysis.
Due to space limitations, errors for are not included here; however, a complete table with errors is available in our dataset. In this study, we select QSOs as foreground pairs for the main sample, as they provide a broad redshift distribution of foreground objects, facilitating a more comprehensive examination of evolutionary effects. The equivalent width results for ELGs and LRGs are included in Appendix A, and the corresponding data, including the stacked spectra and absorption measurements, can be accessed on the same dataset. The impact parameter bins, ranging from 0.02 to 6.5 Mpc, encompass the CGM and the gas that connects to the large-scale structure of the universe, making them suitable for conducting a thorough survey of the gas associated with QSOs.
@@ 3.1 Evolution of with redshift and enrichment rates
We plot the data from Table LABEL:tab1 that are not marked with an asterisk and are sufficiently removed from the blue end of the stacked spectrum in Figure 2. This selection criterion is consistently applied throughout all figures in this paper. The upper panels display the equivalent widths () using different marker shapes to represent each element, with colors indicating their respective rest wavelengths. This marking scheme is consistent across all figures in this work, except for Figure 6, which focuses on the ratio of two lines. The lower panels of Figure 2 present the normalized equivalent widths () as a function of .
A consistent negative correlation is observed across all ions, where systematically decreases as redshift increases. The lower panels illustrate this trend more distinctly. The normalization of accounts for differences in the initial values, facilitating clearer comparisons of trends across various radial bins and redshifts. The observed decrease in with redshift is evident in all radial bins with slopes of (0.8 Mpc), (2.5 Mpc), and (5.1 Mpc).
We also fit the evolution separately for ions tracing cold gas (e.g., C i, Mg i), cool gas (e.g., Fe ii, Mg ii), and warm gas (e.g., C iv, S iv), as well as for -process elements (e.g., Mg, O, Si) and Fe-peak elements (e.g., Fe, Ni, Mn). Variations in the nucleosynthetic origins of these elements may lead to distinct evolutionary trends over cosmic time. In general, -elements show slightly weaker evolution at compared to Fe-peak elements, while absorption lines tracing warmer gas exhibit somewhat stronger evolution. Nevertheless, given the relatively large uncertainties in the fitted slopes, these comparisons should be interpreted with caution.
In Figure 3, we present the fitted slopes for each absorption line in the vs. plot (as shown in Figure 2). The evolution of different elements and ions shows significant variability. For example, among all the absorption lines, Mg ii shows the weakest evolution, indicating little to no change over time, which aligns with trends noted by Lan (2020), Anand et al. (2021) and Chen et al. (2025). The overall structure of Mg ii-bearing halos was established early in the galaxy assembly process and have evolved relatively little since then (Matejek & Simcoe, 2012).
@@
We also perform the linear fittings to the profiles of these absorption lines as shown in the right panel of Figure 4. The results indicates that the absorption intensity decreases with , which aligns with previous findings (e.g., Zhu & Ménard, 2013a; Lan & Mo, 2018; Méndez-Hernández et al., 2022; Wu et al., 2025). Mg ii absorption strength here is slightly lower than Zhu et al. (2014) for LRGs at . This discrepancy primarily arises from the higher redshift of the foreground QSOs () in our study. Ly demonstrates a stronger absorption as predicted by Shen et al. (2013) and remains relatively stable on large scales. Even after accounting for increased model complexity, the halo model still provides a better overall fit (with a smaller ) than the linear model. In Figure 4, the data points at 5 Mpc appear higher than both models. This may be due to the effect of QSO proximity zones (Zheng et al., 2015; Onorato et al., 2025), beyond which the ionization influence of the host QSO diminishes. This reduction in ionization could lead to an increase in the fraction of low-ionization species.
The fitted slopes from the linear fittings for each ion are displayed in Figure 5. The observed differences in could be attributed to the varying physical processes that affect the distribution of each ion in the gas surrounding galaxies. For example, high-ionization species such as C iv and Si iv are typically found in hotter, more ionized regions. In contrast, low-ionization species such as Mg ii and Fe ii are generally located in colder, and denser regions. High-ionization species tend to maintain a more extended distribution (with smaller ), as corroborated by both simulations (Shen et al., 2013; Sanchez et al., 2019; Li et al., 2021; Carr et al., 2025) and observational studies (Tumlinson et al., 2011; Prochaska et al., 2011b; Méndez-Hernández et al., 2022). Conversely, low-ionization species Mg ii and Fe ii show stronger dependence on impact factor, with values of (Fe ii ), (Fe ii ), (Fe ii ), (Fe ii ), and (Mg ii ). Additionally, in Figure 5, we present the slope of the projected two-point correlation function for galaxies within the 1–5 Mpc range, marked by different colors: Zehavi et al. (2004) for (blue), Nuza et al. (2013) for (green), and Durkalec et al. (2018) for (orange).
@@
The ratio is shown as a function of (left panel) and the impact parameter (right panel) in Figure 6. Different colors correspond to various Fe ii absorption lines indicated in the legend. The results indicate that as the redshift increases, the ratio decreases following a trend proportional to , consistent with Dey et al. (2015). This trend is consistently observed within each individual impact parameter bin. In contrast, the variation of with distance is not significant within the 1- confidence level.
Mg, being an -element, is predominantly produced by core-collapse supernovae (CCSN). In contrast, iron (Fe) is primarily generated by Type Ia supernovae (SNIa), which originate from the explosion of accreting white dwarfs. The ratio of these elements should reflect the relative frequencies of CCSN and SNIa, as well as the time lag between their productions. In studying the stellar absorption lines, Conroy et al. (2014) found significant enhancements in massive elliptical galaxies formed in the early universe, attributed to the intense bursts of star formation that occurred at high redshift (Li et al., 2025).